# A Comprehensive Survey for Hyperspectral Image Classification: The Evolution from Conventional to Transformers

### v3

@@ 11 Transformers for HSC
Considering these limitations, various potential solutions have been proposed in the literature, including the integration of SSTs with other architectures. For instance, a hybrid model could leverage the strengths of both approaches by combining the global context understanding of SSTs with the local feature extraction capabilities of CNNs [4, 313]. However, a potential drawback of hybrid Transformers is the increased complexity in model architecture. Integrating different Transformer variants or combining Transformers with other architectures might introduce additional intricacy, rendering the model more challenging to interpret and potentially demanding more computational resources for training and inference.
@@
Moreover, there exists a realm for exploration in fine-tuning the attention mechanism, where endeavors concentrate on sparse attention patterns or alternative attention mechanisms aimed at tackling the computational complexity challenges [314, 315, 302]. Capturing long-range dependencies and integrating global and local features in Transformers entails the challenge of finding a delicate equilibrium between model complexity and computational efficiency. Although optimizing attention patterns can elevate performance, it may concurrently heighten computational requirements, rendering the model more demanding on resources and possibly constraining its utility in scenarios with limited resources.
@@
Expanding the capabilities of SSTs to accommodate multi-modal inputs emerges as a promising avenue for further research [316]. This extension has the potential to widen their applicability and augment their ability to comprehend intricate relationships within diverse data types [92, 317, 318]. However, integrating information from various modalities necessitates thoughtful consideration of feature representations and alignment. The model might encounter difficulties in effectively learning meaningful cross-modal relationships, and crafting architectures capable of efficiently fusing and processing diverse data sources remains an ongoing research challenge. Additionally, gathering and annotating large-scale multi-modal datasets for training can be resource-intensive. Lastly, developing efficient tokenization strategies is crucial, aiming to capture fine-grained details while alleviating the quadratic scaling issue. Such strategies could significantly boost the performance of SSTs in tasks requiring high spatial resolution. In summary, while both ST and SST are Transformer-based architectures tailored for image classification tasks, they diverge in their approach to handling image data. Here are some of the major differences between ST and SST:
@@
11.1 Benefits and Challenges of Transformers in HS Analysis
@@
11.2 Emerging Trends and Advancements
As HS technology continues to evolve, several emerging trends and advancements in HSC have garnered attention within the remote sensing and image-processing communities. These trends reflect ongoing efforts to enhance the accuracy, efficiency, and applicability to HS data analysis. The integration of DL and Transformer-based models has revolutionized HSC. DL architectures excel at automatically learning hierarchical and complex features from HS data, improving classification accuracy. Transformers, with their ability to capture global dependencies, are being explored to address challenges related to long-range spectral interactions. In the recent past, several trends have emerged, aiming to enhance the overall generalization performance of DL models, and a few of them are listed in the following.
11.2.1 Domain Adaptation and Transfer Learning
@@
In DNN-based HSC, the model undergoes hierarchical feature learning, with lower layers typically extracting generic features when trained on diverse images. Consequently, the features acquired by these lower layers can be transferred to train a new classifier for the target dataset. For example, Yang et al. [319] employed a two-branch spectral-spatial CNN model, initially trained on a substantial amount of data from other HSIs. The lower layers of this pre-trained model were then applied to the target network, ensuring robust classification of the target HSI. To capture target-specific features, the higher layers of the target network were randomly initialized, and the entire network underwent fine-tuning using limited labeled instances from the target HSI. Similarly, Windrim et al. [320] introduced an effective method for pre-training and fine-tuning a CNN network, making it adaptable for the classification of new HSIs. The study conducted by Liu et al. [321] incorporated both data augmentation and transfer learning approaches to address the shortage of training data, aiming to enhance the performance of HSC.
@@
As discussed above, data in the source and target domains may differ in various aspects. For instance, in the context of HSIs, the dimensions of two HSIs may vary due to acquisition from different sensors. Addressing such cross-domain variations and facilitating knowledge transfer between them is referred to as heterogeneous transfer learning, and a comprehensive survey of such methods can be found in [322]. Within the literature on HSC, numerous studies have been put forth to bridge the gap and transfer knowledge between two HSIs with varying dimensions and/or distributions. For instance, an effective framework for HSC based on heterogeneous transfer learning was introduced by Lin et al. [323]. This framework demonstrates efficacy with both homogeneous and heterogeneous HSIs. Another approach, proposed by Li et al. [324], utilizes an iterative re-weighting mechanism within a heterogeneous transfer learning framework for HSC. A work proposed by Liu et al. [325] advocates a band selection-based transfer learning approach to pre-train a CNN, maintaining a consistent number of dimensions across various HSIs. Additionally, Lin et al. put forward an unsupervised transfer learning technique designed to classify entirely unknown target HSIs. Furthermore, Pires et al. [326] demonstrate that networks trained on natural images can enhance the performance of transfer learning for remote sensing data classification compared to networks trained from scratch using smaller HS datasets.
11.2.2 Data Augmentation
@@
Augmentation is a proven and effective technique in HSC to address the challenge of limited training samples. It involves generating new samples from the original training data without the need for additional labeling costs. Data augmentation approaches can be broadly categorized into two main strategies: Data Wrapping: This strategy focuses on encoding invariances such as translational, size, viewpoint, and illumination variations by applying geometric and color-based transformations while preserving the original labels. These transformations help create new samples that exhibit similar characteristics to the original data but with variations that can improve the model’s robustness and generalization. Examples of data wrapping techniques include geometric transformations like rotation, scaling, and flipping, as well as color-based transformations like brightness and contrast adjustments. Oversampling: Oversampling-based augmentation methods aim to inflate the training data by generating synthetic samples based on the original data distribution [327]. These techniques help address class imbalance issues and provide the model with a more balanced representation of different classes. Oversampling techniques include mixture-based instance generation, where new samples are created by blending existing samples from the same class, and feature space augmentations, where new samples are generated by perturbing the feature vectors of existing samples.
@@
In the realm of HSC literature, various frameworks utilizing data augmentation techniques have been implemented to enhance classification performance and mitigate potential overfitting issues, often stemming from limited training data availability. For example, Yu et al. [328] improved training data by applying three augmentation operations (flip, rotate, and translation), leveraging this augmented dataset to train CNN for HSC. The study by Cao et al. [329] assessed the classification performance of a combination of CNN and Active Learning (AL), both with and without data augmentation techniques, demonstrating higher accuracy when employing data augmentation. Similarly, in another comparative study [330], a data augmentation-based CNN exhibited a 10% increase in HSC accuracy compared to a PCA-based CNN model.
@@
The aforementioned approaches employ offline data augmentation methods, which involve increasing the training data by generating new instances during or before the model training process. A recent innovative data augmentation framework for HSI is introduced in [331]. In contrast to expanding the training data, this framework generates samples at test time. A DNN trained on the original training data, coupled with a voting scheme, is utilized to determine the final class label. To enhance the generalization capability of DNN models, [331] also puts forth two rapid data augmentation techniques for high-quality data synthesis. Additionally, a similar online data augmentation strategy based on PCA is proposed in [332]. This strategy synthesizes new instances during the inference phase rather than during the training phase.
11.2.3 Generative Adversarial Networks
@@
The Generative Adversarial Network (GAN), introduced by Goodfellow et al. [333], consists of two neural networks: a generator and a discriminator. GANs can learn and replicate samples by leveraging intricate details within the data distribution. In the context of Semi-Supervised Learning-based HSC, Zhan et al. [334] presented a GAN approach that relies on spectral features. In a similar vein, He et al. [335] introduced a GAN-based framework for spectral-spatial HSC. Likewise, Zhu et al. [336] devised 1D-GAN and 3D-GAN architectures, based on CNN, aimed at enhancing classification performance. Zhan et al. [337] employed a customized 1D-GAN to generate spectral features, utilized by a CNN for subsequent feature extraction, and concluded the process with majority voting in HSC. Recently, Feng et al. [338] presented a Spatial-Spectral Multi-Class GAN (MSGAN) employing two generators to produce spatial and spectral information through multiple adversarial objectives. In addressing data imbalance issues for HSC, Zhong et al. [339] proposed a semi-supervised model combining GAN with Conditional Random Fields (CRFs).
@@
Similarly, Wang et al. [340] explored the Caps-TripleGAN model, which efficiently generates new samples through a 1D structure Triple GAN (TripleGAN) and classifies the generated HSI samples using the capsule network (CapsNet). Additionally, Xue et al. [341] suggested the adoption of a 3D CNN-based generator network and a 3D deep residual network-based discriminator network for HSC. To learn high-level contextual features, the combination of a capsule network and a Convolutional LSTM (ConvLSTM) based discriminator model was introduced by Wang et al. [342] in the context of HSC. The study conducted by Alipour et al. [343] aimed to overcome the limitation of scarce training examples by employing a GAN model. In this approach, the discriminator’s performance is enhanced through an auxiliary classifier, generating more structurally coherent virtual training samples. Additionally, Roy et al. [344] introduced a generative adversarial minority oversampling-based technique to augment model performance. This approach specifically addresses the persistent challenge of class-wise data imbalance in HSC.
11.2.4 Active Learning (AL)
@@
Given the cost and effort associated with acquiring labeled HS data, AL strategies are becoming more prevalent. These techniques involve selecting the most informative samples for annotation, iteratively improving the model’s performance with minimal labeled data. AL helps in optimizing the learning process, making HSC more efficient and cost-effective. The process of selecting the most informative samples is performed to ensure that the chosen samples are both informative and representative of the overall input distribution, ultimately enhancing accuracy. Depending on the criteria for adding new instances to the training set, AL frameworks can be categorized as either stream-based or pool-based. In the stream-based selection, one instance is drawn at a time from an existing set of unlabeled samples, and the model determines whether to label it or not based on its perceived usefulness. Conversely, in the pool-based strategy, samples are queried from a pool or subset of unlabeled data. This selection is based on ranking scores computed from various measures to assess the sample’s usefulness. The study by Ganti et al. [345] observed that stream-based selection yields lower learning rates compared to pool-based selection, primarily due to the former’s tendency to query additional instances. In pool-based selection, emphasizing diversity within the pool of samples is crucial to prevent redundancy. Typically, the selection or querying of the most valuable samples centers around three main aspects: heterogeneity behavior, model performance, and representativeness of samples.
@@
Performance-based Selection: These methods take into account the impact of adding queried samples on the model’s performance. They aim to optimize the model’s performance by minimizing variance and error. Performance-based sampling can be broadly categorized into two types i.e., Expected Error Reduction [346] and Expected Variance Reduction [347].
@@
Representativeness-based selection: Representative sampling tends to query instances that are representative of the overall input distribution, thus steering clear of outliers and unrepresentative samples. These methods assign a higher importance to the denser input region during the querying process. Representativeness sampling approaches, such as density-weighted techniques like information density, consider both the representativeness of samples and heterogeneity behavior. These models are often referred to as hybrid models [346].
@@
AL has gained substantial prominence in recent times within HSC. Liu et al. [348] introduced a feature-driven AL framework, aiming to establish a well-constructed feature space for HSC. Zhang et al. [349] proposed a semi-supervised AL method based on Random Forest, leveraging spectral-spatial features to formulate a query function. This function is employed to select the most informative samples as target candidates for the training set. Guo et al. [350] introduced a framework that integrates the spectral and spatial features of superpixels. Likewise, Xue et al. [351] incorporated neighborhood and superpixel information to augment the uncertainty of queried samples. In a recent study, Bhardwaj et al. [352] leveraged attribute profiles to integrate spatial information within an AL-based HSC framework.
@@
The diversity of samples becomes crucial in batch-mode techniques to prevent redundancy. A multi-criteria batch-mode method introduced by Patra et al. [353] defines a novel query function based on diversity, uncertainty, and cluster assumption measures. These criteria leverage the properties of KNN, SVM, and K-means clustering, respectively. Genetic algorithms are subsequently employed to choose the most effective batch of samples. Similarly, Zhang et al. [354] proposed a regularized multi-metric batch-mode framework for HSC, exploiting various features of HSI. A Multiview AL (MVAL) framework was introduced by Xu et al. [355]. This framework analyzes objects from various perspectives and gauges sample informativeness through multiview Intensity-based query criteria. Similarly, Pradhan et al. [356] embraced the concept of multiview learning, employing the Fisher Discriminant Ratio to generate multiple views. In another study, Zhang et al. [357] put forth an innovative adaptive MVAL framework for HSC, simultaneously leveraging spatial and spectral features in each view. More recently, Li et al. [358] proposed an MVAL technique utilizing pixel-level, subpixel-level, and superpixel-level details to generate multiple views for HSC. Furthermore, the proposed method exploits joint posterior probability estimation and dissimilarities among multiple views to query representative samples.
@@
Sun et al. [359] combined an AE with an AL technique, while Liu et al. [360] introduced a DBN-based AL framework for HSC. Similarly, Haut et al. [361] married a Bayesian CNN with the AL paradigm for spectral-spatial HSC. More recently, Cao et al. [329] proposed a CNN-based AL framework to leverage unlabeled samples for HSC effectively. Several studies have incorporated AL with transfer learning for HSC [50]. For instance, Lin et al. [362] introduced a framework that identifies salient samples and utilizes high-level features to establish correlations between source and target domain data. Another approach, suggested by Deng et al. [363], presented an Active Transfer Learning technique based on Stacked Sparse AE. This technique jointly utilizes both spectral and spatial features for HSC. Additionally, Deng et al. [364] combined domain adaptation and AL methods, incorporating multiple kernels for HSC.
@@
AL in HSC introduces sophisticated frameworks aimed at improving the generalization capabilities of models. For example, Ahmad et al. [365] presented a fuzziness-based to enhance the generalization performance of both discriminative and generative classifiers. This method calculates the fuzziness-based distance for each instance and the estimated class boundary. Instances with higher fuzziness values and smaller distances from class boundaries are selected as candidates for the training set. Recently, a non-randomized spectral-spatial framework for multiclass HSC was introduced in [366]. This framework combines the spatial prior fuzziness approach with Multinomial Logistic Regression via a Splitting and Augmented Lagrangian classifier. The authors conducted a comprehensive comparison of the proposed framework with state-of-the-art sample selection methods and diverse classifiers.
11.2.5 Graph-based Approaches
Graph-based methods are gaining popularity for HSC, especially in scenarios where spatial-spectral relationships are crucial. GCNs and graph attention networks (GATs) enable the modeling of intricate interactions between neighboring pixels, capturing both spectral and spatial information. These methods have shown promise in improving classification accuracy, particularly in urban and complex landscapes.
11.2.6 Explainable AI for HS Analysis
@@
11.2.7 Integration of Multimodal Data
Combining HS data with other imaging modalities, such as Light Detection and Ranging (LiDAR) or Synthetic Aperture Radar (SAR), is an emerging trend. The fusion of multimodal data sources enhances the overall understanding of the scene [367], providing complementary information for improved classification accuracy and discrimination of complex land cover classes.
11.2.8 Edge Computing for Real-time Processing
@@
11.2.9 Development of Benchmark Datasets
@@
In short, the field of HSC is witnessing rapid advancements driven by innovations in DL, domain adaptation, graph-based methods, AL, explainable AI, multimodal data integration, edge computing, and the establishment of benchmark datasets [368]. These trends collectively contribute to the refinement of HS data analysis techniques and expand the applicability of HSI in various domains. As technology continues to progress, further breakthroughs will likely shape the future landscape of HSC.

### v4

@@ 11 Transformers for HSC
Considering these limitations, various potential solutions have been proposed in the literature, including the integration of SSTs with other architectures. For instance, a hybrid model could leverage the strengths of both approaches by combining the global context understanding of SSTs with the local feature extraction capabilities of CNNs (Cao et al., 2024b, ; Li et al., 2024f, ). However, a potential drawback of hybrid Transformers is the increased complexity in model architecture. Integrating different Transformer variants or combining Transformers with other architectures might introduce additional intricacy, rendering the model more challenging to interpret and potentially demanding more computational resources for training and inference.
@@
Moreover, there exists a realm for exploration in fine-tuning the attention mechanism, where endeavors concentrate on sparse attention patterns or alternative attention mechanisms aimed at tackling the computational complexity challenges (Shu et al.,, 2024; Ma et al., 2024a, ; Ma et al., 2024b, ). Capturing long-range dependencies and integrating global and local features in Transformers entails the challenge of finding a delicate equilibrium between model complexity and computational efficiency. Although optimizing attention patterns can elevate performance, it may concurrently heighten computational requirements, rendering the model more demanding on resources and possibly constraining its utility in scenarios with limited resources.
@@
Expanding the capabilities of SSTs to accommodate multi-modal inputs emerges as a promising avenue for further research (Zhang et al., 2024d, ). This extension has the potential to widen their applicability and augment their ability to comprehend intricate relationships within diverse data types (Li et al., 2024a, ; Miao et al.,, 2024; Liu et al.,, 2021). However, integrating information from various modalities necessitates thoughtful consideration of feature representations and alignment. The model might encounter difficulties in effectively learning meaningful cross-modal relationships, and crafting architectures capable of efficiently fusing and processing diverse data sources remains an ongoing research challenge. Additionally, gathering and annotating large-scale multi-modal datasets for training can be resource-intensive. Lastly, developing efficient tokenization strategies is crucial, aiming to capture fine-grained details while alleviating the quadratic scaling issue. Such strategies could significantly boost the performance of SSTs in tasks requiring high spatial resolution. In summary, while both ST and SST are Transformer-based architectures tailored for image classification tasks, they diverge in their approach to handling image data. Here are some of the major differences between ST and SST:
@@
11.a Benefits and Challenges of Transformers
@@
11.b Emerging Trends and Advancements
As HS technology continues to evolve, several emerging trends and advancements in HSC have garnered attention within the remote sensing and image-processing communities. These trends reflect ongoing efforts to enhance the accuracy, efficiency, and applicability of HS data analysis. The integration of DL and Transformer-based models has revolutionized HSC. DL architectures excel at automatically learning hierarchical and complex features from HS data, improving classification accuracy. Transformers, with their ability to capture global dependencies, are being explored to address challenges related to long-range spectral interactions. In the recent past, several trends have emerged, aiming to enhance the overall generalization performance of DL models, and a few of them are listed in the following.
11.b.i Domain Adaptation and Transfer Learning
@@
In DNN-based HSC, the model undergoes hierarchical feature learning, with lower layers typically extracting generic features when trained on diverse images. Consequently, the features acquired by these lower layers can be transferred to train a new classifier for the target dataset. For example, Yang et al. (Yang et al.,, 2017) employed a two-branch spectral-spatial CNN model, initially trained on a substantial amount of data from other HSIs. The lower layers of this pre-trained model were then applied to the target network, ensuring robust classification of the target HSI. To capture target-specific features, the higher layers of the target network were randomly initialized, and the entire network underwent fine-tuning using limited labeled instances from the target HSI. Similarly, Windrim et al. (Windrim et al.,, 2018) introduced an effective method for pre-training and fine-tuning a CNN network, making it adaptable for the classification of new HSIs. The study conducted by Liu et al. (Liu et al.,, 2018) incorporated both data augmentation and transfer learning approaches to address the shortage of training data, aiming to enhance the performance of HSC.
@@
As discussed above, data in the source and target domains may differ in various aspects. For instance, in the context of HSIs, the dimensions of two HSIs may vary due to acquisition from different sensors. Addressing such cross-domain variations and facilitating knowledge transfer between them is referred to as heterogeneous transfer learning, and a comprehensive survey of such methods can be found in (Day and Khoshgoftaar,, 2017). Within the literature on HSC, numerous studies have been put forth to bridge the gap and transfer knowledge between two HSIs with varying dimensions and/or distributions. For instance, an effective framework for HSC based on heterogeneous transfer learning was introduced by Lin et al. (Li et al., 2017a, ). This framework demonstrates efficacy with both homogeneous and heterogeneous HSIs. Another approach, proposed by Li et al. (Liu and Xiao,, 2020), utilizes an iterative re-weighting mechanism within a heterogeneous transfer learning framework for HSC. A work proposed by Liu et al. (Lin et al.,, 2017) advocates a band selection-based transfer learning approach to pre-train a CNN, maintaining a consistent number of dimensions across various HSIs. Additionally, Lin et al. put forward an unsupervised transfer learning technique designed to classify entirely unknown target HSIs. Furthermore, Pires et al. (Pires de Lima and Marfurt,, 2020) demonstrate that networks trained on natural images can enhance the performance of transfer learning for remote sensing data classification compared to networks trained from scratch using smaller HS datasets.
11.b.ii Data Augmentation
@@
Augmentation is a proven and effective technique in HSC to address the challenge of limited training samples. It involves generating new samples from the original training data without the need for additional labeling costs. Data augmentation approaches can be broadly categorized into two main strategies: Data Wrapping: This strategy focuses on encoding invariances such as translational, size, viewpoint, and illumination variations by applying geometric and color-based transformations while preserving the original labels. These transformations help create new samples that exhibit similar characteristics to the original data but with variations that can improve the model’s robustness and generalization. Examples of data wrapping techniques include geometric transformations like rotation, scaling, and flipping, as well as color-based transformations like brightness and contrast adjustments. Oversampling: Oversampling-based augmentation methods aim to inflate the training data by generating synthetic samples based on the original data distribution (Shorten and Khoshgoftaar,, 2019). These techniques help address class imbalance issues and provide the model with a more balanced representation of different classes. Oversampling techniques include mixture-based instance generation, where new samples are created by blending existing samples from the same class, and feature space augmentations, where new samples are generated by perturbing the feature vectors of existing samples.
@@
In the realm of HSC literature, various frameworks utilizing data augmentation techniques have been implemented to enhance classification performance and mitigate potential overfitting issues, often stemming from limited training data availability. For example, Yu et al. (Yu et al.,, 2017) improved training data by applying three augmentation operations (flip, rotate, and translation), leveraging this augmented dataset to train CNN for HSC. Similarly, in another comparative study (Rochac et al.,, 2019), a data augmentation-based CNN exhibited a 10% increase in HSC accuracy compared to a PCA-based CNN model.
@@
The aforementioned approaches employ offline data augmentation methods, which involve increasing the training data by generating new instances during or before the model training process. A recent innovative data augmentation framework for HSI is introduced in (Nalepa et al.,, 2020). In contrast to expanding the training data, this framework generates samples at test time. A DNN trained on the original training data, coupled with a voting scheme, is utilized to determine the final class label. To enhance the generalization capability of DNN models, (Nalepa et al.,, 2020) also puts forth two rapid data augmentation techniques for high-quality data synthesis. Additionally, a similar online data augmentation strategy based on PCA is proposed in (Nalepa et al.,, 2019). This strategy synthesizes new instances during the inference phase rather than during the training phase.
11.b.iii Generative Adversarial Networks
@@
The Generative Adversarial Network (GAN), introduced by Goodfellow et al. (Goodfellow et al.,, 2014), consists of two neural networks: a generator and a discriminator. GANs can learn and replicate samples by leveraging intricate details within the data distribution. In the context of Semi-Supervised Learning-based HSC, Zhan et al. (Zhan et al.,, 2017) presented a GAN approach that relies on spectral features. In a similar vein, He et al. He et al., 2017b introduced a GAN-based framework for spectral-spatial HSC. Likewise, Zhu et al. (Zhu et al., 2018b, ) devised 1D-GAN and 3D-GAN architectures, based on CNN, aimed at enhancing classification performance. Zhan et al. Zhan et al., (2018) employed a customized 1D-GAN to generate spectral features, utilized by a CNN for subsequent feature extraction, and concluded the process with majority voting in HSC. Recently, Feng et al. (Feng et al.,, 2019) presented a Spatial-Spectral Multi-Class GAN (MSGAN) employing two generators to produce spatial and spectral information through multiple adversarial objectives. In addressing data imbalance issues for HSC, Zhong et al. (Zhong et al.,, 2019) proposed a semi-supervised model combining GAN with Conditional Random Fields (CRFs).
@@
Similarly, Wang et al. (Wang et al.,, 2019) explored the Caps-TripleGAN model, which efficiently generates new samples through a 1D structure Triple GAN (TripleGAN) and classifies the generated HSI samples using the capsule network (CapsNet). Additionally, Xue et al. (Xue,, 2019) suggested the adoption of a 3D CNN-based generator network and a 3D deep residual network-based discriminator network for HSC. To learn high-level contextual features, the combination of a capsule network and a Convolutional LSTM (ConvLSTM) based discriminator model was introduced by Wang et al. (Wang et al.,, 2020) in the context of HSC. The study conducted by Alipour et al. (Alipour-Fard and Arefi,, 2020) aimed to overcome the limitation of scarce training examples by employing a GAN model. In this approach, the discriminator’s performance is enhanced through an auxiliary classifier, generating more structurally coherent virtual training samples. Additionally, Roy et al. (Roy et al., 2021a, ) introduced a generative adversarial minority oversampling-based technique to augment model performance. This approach specifically addresses the persistent challenge of class-wise data imbalance in HSC.
11.b.iv Active Learning (AL)
@@
Given the cost and effort associated with acquiring labeled HS data, AL strategies are becoming more prevalent. These techniques involve selecting the most informative samples for annotation, iteratively improving the model’s performance with minimal labeled data. AL helps in optimizing the learning process, making HSC more efficient and cost-effective. The process of selecting the most informative samples is performed to ensure that the chosen samples are both informative and representative of the overall input distribution, ultimately enhancing accuracy. Depending on the criteria for adding new instances to the training set, AL frameworks can be categorized as either stream-based or pool-based. In the stream-based selection, one instance is drawn at a time from an existing set of unlabeled samples, and the model determines whether to label it or not based on its perceived usefulness. Conversely, in the pool-based strategy, samples are queried from a pool or subset of unlabeled data. This selection is based on ranking scores computed from various measures to assess the sample’s usefulness. The study by Ganti et al. (Ganti and Gray,, 2012) observed that stream-based selection yields lower learning rates compared to pool-based selection, primarily due to the former’s tendency to query additional instances. In pool-based selection, emphasizing diversity within the pool of samples is crucial to prevent redundancy. Typically, the selection or querying of the most valuable samples centers around three main aspects: heterogeneity behavior, model performance, and representativeness of samples.
@@
Performance-based Selection: These methods take into account the impact of adding queried samples on the model’s performance. They aim to optimize the model’s performance by minimizing variance and error. Performance-based sampling can be broadly categorized into two types i.e., Expected Error Reduction (Aggarwal et al.,, 2014) and Expected Variance Reduction (Settles,, 2009).
@@
Representativeness-based selection: Representative sampling tends to query instances that are representative of the overall input distribution, thus steering clear of outliers and unrepresentative samples. These methods assign a higher importance to the denser input region during the querying process. Representativeness sampling approaches, such as density-weighted techniques like information density, consider both the representativeness of samples and heterogeneity behavior. These models are often referred to as hybrid models (Aggarwal et al.,, 2014).
@@
AL has gained substantial prominence in recent times within HSC. Liu et al. (Liu et al., 2017a, ) introduced a feature-driven AL framework, aiming to establish a well-constructed feature space for HSC. Zhang et al. (Zhang et al., 2019b, ) proposed a semi-supervised AL method based on Random Forest, leveraging spectral-spatial features to formulate a query function. This function is employed to select the most informative samples as target candidates for the training set. Guo et al. (Guo et al.,, 2016) introduced a framework that integrates the spectral and spatial features of superpixels. Likewise, Xue et al. (Xue et al.,, 2018) incorporated neighborhood and superpixel information to augment the uncertainty of queried samples. In a recent study, Bhardwaj et al. (Bhardwaj et al.,, 2020) leveraged attribute profiles to integrate spatial information within an AL-based HSC framework.
@@
The diversity of samples becomes crucial in batch-mode techniques to prevent redundancy. A multi-criteria batch-mode method introduced by Patra et al. (Patra et al.,, 2017) defines a novel query function based on diversity, uncertainty, and cluster assumption measures. These criteria leverage the properties of KNN, SVM, and K-means clustering, respectively. Genetic algorithms are subsequently employed to choose the most effective batch of samples. Similarly, Zhang et al. (Zhang and Crawford,, 2017) proposed a regularized multi-metric batch-mode framework for HSC, exploiting various features of HSI. A Multiview AL (MVAL) framework was introduced by Xu et al. (Xu et al.,, 2017). This framework analyzes objects from various perspectives and gauges sample informativeness through multiview Intensity-based query criteria. Similarly, Pradhan et al. (Pradhan et al.,, 2018) embraced the concept of multiview learning, employing the Fisher Discriminant Ratio to generate multiple views. In another study, Zhang et al. (Zhang et al., 2019c, ) put forth an innovative adaptive MVAL framework for HSC, simultaneously leveraging spatial and spectral features in each view. More recently, Li et al. (Li et al.,, 2020) proposed an MVAL technique utilizing pixel-level, subpixel-level, and superpixel-level details to generate multiple views for HSC. Furthermore, the proposed method exploits joint posterior probability estimation and dissimilarities among multiple views to query representative samples.
@@
Sun et al. (Sun et al.,, 2016) combined an AE with an AL technique, while Liu et al. (Liu et al.,, 2016) introduced a DBN-based AL framework for HSC. Similarly, Haut et al. (Haut et al.,, 2018) married a Bayesian CNN with the AL paradigm for spectral-spatial HSC. More recently, Cao et al. (Cao et al.,, 2020) proposed a CNN-based AL framework to leverage unlabeled samples for HSC effectively. Several studies have incorporated AL with transfer learning for HSC (Ahmad et al., 2022a, ). For instance, Lin et al. (Lin et al.,, 2018) introduced a framework that identifies salient samples and utilizes high-level features to establish correlations between source and target domain data. Another approach, suggested by Deng et al. (Deng et al., 2018b, ), presented an Active Transfer Learning technique based on Stacked Sparse AE. This technique jointly utilizes both spectral and spatial features for HSC. Additionally, Deng et al. (Deng et al., 2018a, ) combined domain adaptation and AL methods, incorporating multiple kernels for HSC.
@@
AL in HSC introduces sophisticated frameworks aimed at improving the generalization capabilities of models. For example, Ahmad et al. (Ahmad et al.,, 2018) presented a fuzziness-based to enhance the generalization performance of both discriminative and generative classifiers. This method calculates the fuzziness-based distance for each instance and the estimated class boundary. Instances with higher fuzziness values and smaller distances from class boundaries are selected as candidates for the training set. Recently, a non-randomized spectral-spatial framework for multiclass HSC was introduced in (Ahmad et al., 2020b, ). This framework combines the spatial prior fuzziness approach with Multinomial Logistic Regression via a Splitting and Augmented Lagrangian classifier. The authors conducted a comprehensive comparison of the proposed framework with state-of-the-art sample selection methods and diverse classifiers.
11.b.v Explainable AI for HS Analysis
@@
11.b.vi Integration of Multimodal Data
Combining HS data with other imaging modalities, such as Light Detection and Ranging (LiDAR) or Synthetic Aperture Radar (SAR), is an emerging trend. The fusion of multimodal data sources enhances the overall understanding of the scene (Hong et al., 2023a, ), providing complementary information for improved classification accuracy and discrimination of complex land cover classes.
In short, multimodal data refers to the combination of various types of data (e.g., spectral, spatial, textual, and contextual) that can provide complementary information for improving the accuracy and robustness of HSC. In the context of HSC, multimodal data can encompass several modalities, including: Spectral Data: Captured through HS sensors, this modality provides continuous spectral profiles for each pixel, enabling the identification of unique spectral signatures associated with different materials. Spatial Data: Information about the spatial arrangement of pixels, including texture and contextual relationships, can significantly enhance classification accuracy. Spatial information can be derived from multi-resolution imagery, such as RGB or multispectral images. Textual Data: Descriptive textual information about the scene, such as land cover type, can be integrated into the classification framework. This data may come from ancillary sources like metadata, environmental reports, or expert annotations. Contextual Data: Additional contextual features, such as elevation, climate conditions, or geographical information, can provide valuable insights into the classification process, especially in heterogeneous landscapes.
Integrating multimodal data offers several advantages for HSC: Enhanced Discriminative Power: By combining spectral and spatial information, classifiers can leverage the complementary strengths of each modality, leading to improved separability of classes. Robustness to Noise: Multimodal integration can mitigate the effects of noise present in HS data, as different modalities can corroborate each other’s information and reduce uncertainty in the classification output. Improved Generalization: Utilizing diverse data sources helps models generalize better across varying conditions, making them less prone to overfitting on training data. Better Representation of Complex Scenes: Multimodal approaches can capture the complexities of real-world environments more effectively, enabling better classification of mixed or ambiguous land cover types.
Several techniques exist for fusing multimodal data in HSC for instance, Early Fusion: In this approach, multimodal data is concatenated at the feature level before classification. For example, spectral features from HS data can be combined with texture features extracted from spatial data. Late Fusion: This method involves separately classifying each modality and then combining the resulting classification maps. Techniques such as voting schemes or probability averaging can be applied to merge the outputs. Hybrid Fusion: A combination of early and late fusion strategies, hybrid approaches aim to leverage the strengths of both methods. For instance, features can be extracted from each modality and then combined with late-stage classification outputs. Deep Learning Approaches: Recent advancements in DL have enabled the development of multimodal neural networks capable of learning joint representations from different modalities. For instance, CNNs can process spatial data, while RNNs can handle sequential textual data. Feature-Level Fusion: Techniques like early fusion concatenate features from multiple modalities at the input stage, allowing for a more holistic representation of the data, while late fusion combines predictions from separate models trained on individual modalities. Attention Mechanisms: Recent approaches incorporate attention mechanisms that dynamically weigh the importance of various modalities based on the classification task, enhancing the model’s ability to focus on relevant features. Graph-Based Methods: Innovations in graph-based methods allow for the modeling of relationships between different data modalities, which can improve the robustness of classification results in complex environments.
Despite its advantages, multimodal data integration presents several challenges: Data Alignment: Ensuring spatial and spectral data are aligned correctly is crucial. Misalignment can lead to inaccurate classifications and degraded performance. Dimensionality Management: The integration of multiple data sources can result in significantly high dimensionality, leading to computational challenges and the curse of dimensionality. Effective dimensionality reduction techniques, such as PCA or t-SNE, may be necessary. Model Complexity: Incorporating multiple modalities often requires more sophisticated models, which can increase the complexity of training and inference processes. Data Availability: In some cases, obtaining multimodal data can be challenging, especially if certain modalities are expensive or difficult to capture.
11.b.vii Edge Computing for Real-time Processing
@@
11.b.viii Development of Benchmark Datasets
@@
In short, the field of HSC is witnessing rapid advancements driven by innovations in DL, domain adaptation, graph-based methods, AL, explainable AI, multimodal data integration, edge computing, and the establishment of benchmark datasets (Hong et al., 2023b, ). These trends collectively contribute to the refinement of HS data analysis techniques and expand the applicability of HSI in various domains. As technology continues to progress, further breakthroughs will likely shape the future landscape of HSC.